40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1663     Entry time: Fri Jul 8 22:09:33 2016
Author: Antonio 
Type: DailyProgress 
Category: optic 
Subject: North path construction progress 2 

North Path

Summary
Following entry ID 1649, additional work along the North Path has been done

Conclusion
We have the light into the North cavity mode mateched and aligned to TEM00 
with a visibility {(Vis  = Vmax-Vmin) / Vmax} of ~71%.

Next Step
If FOR THE MOMENT we accept this mode matching I will hook up the electronics
and work on the North cavity locking, leaving modematching improvments for later.




Things done
1. PMC has been placed in the path and aligned. Visibility was about ~90%;
2. PMC has been removed;
3. A PLCX-51.5-UV lens has been cleaned and placed in the optical path soon after the PMC;
4. A resonant EOM (14.5MHz) for PDH sidebandshas been placed and aligned:

  • Power at the input was ~3mW, at the output ~2.95mW

5. Two steering mirrors have been cleaned and placed in the path;
6. Two mode matching lenses have been placed:

  • A PLCX-51.5-UV lens has been placed in a location that few inches above the nominal
    found in the design (ID 1661) ;
     
  • A PLCX-103.0-UV lens has been placed in its nominal location;

7.  In between the two lenses a FI (IO-5-1064-HP) has been aligned:

8. The TEM00 has been found by scanning the PZT of the laser with a triangular wave (+-1V) at 10Hz through an HV amplifier set at ~40V;
    Additionally a voltage calibrator has been used to actuate on the crystal temperature of the laser in order to localize the are to be scanned with the PZT.
 

9. A Lambda/2 has been placed after the FI in order to have one linear polirized light:

  • With light at ~45degree I have seen 2 TEM00 resonance not exactly the same;
  • By rotating the waveplate we can have purely one or the other resonance (not sure yet at which angle);

10. A lens and a steering mirror have been placed in the path of the rejected light from FI to monitor the visibility/resonances with the DC output of the photodiodes;
    For now the visibility is at ~71%

11. This is the current setup:

 

Attachment 1: IMG_1095.jpg  1.938 MB  | Hide | Hide all
IMG_1095.jpg
Attachment 3: North_layout_progress_dist_MM_final.pdf  78 kB  | Hide | Hide all
North_layout_progress_dist_MM_final.pdf
Attachment 4: North_layout_progress_dist_MM_final.pdf  78 kB  | Hide | Hide all
North_layout_progress_dist_MM_final.pdf
ELOG V3.1.3-